# Realising VCD for untwisted automorphism groups of RAAGs

### v1

@@ 7.1. A sufficient condition for $M(L)=M(V)$
Therefore $\overline{Q}$ and $\overline{Q^{\prime}}$ are both hugged. The only vertex which dominates $u_{1}$ and is contained in $\overline{Q}$ is $a_{2}$—so the partition(s) which hug $\overline{Q}$ are based at $a_{2}$. In a hugging set for $\overline{Q}$, there must be a partition $\mathcal{P}$ which has side $P\subseteq\overline{Q}$ with $u_{2}\in P$. Since $\{a_{2}^{\pm}\}\subseteq\overline{Q^{\prime}}$, both sides of $\mathcal{P}$ must intersect $\overline{Q^{\prime}}$. Since $u_{2}\in Q^{\prime}\cap P$, we are forced to have $\overline{P}\cap Q^{\prime}=\emptyset$. Hence $P\supseteq\{u_{2},a_{3}^{\pm},b_{3}^{\pm}\}$.
We already said that the elements $\{u_{2}\},\{a_{3}^{\pm},b_{3}^{\pm}\}$ of $\mathcal{C}(a_{2})$ must be on different sides of $\mathcal{R}$. By the previous paragraph, both sides of $\mathcal{R}$ thus have non-empty intersection with $P$. But the elements $\{u_{1}\},\{a_{1}^{\pm},b_{1}^{\pm}\}$ of $\mathcal{C}(a_{2})$ are also on different sides of $\mathcal{R}$—so both sides of $\mathcal{R}$ intersect $\overline{P}$. Hence $\mathcal{P}$ is incompatible with $\mathcal{R}$, which is a contradiction.
Hence, even though $\Delta$ violates Condition 2, it still satisfies the conclusion of Lemma 4.13. Since $\Delta$ is barbed, we can apply Theorem 5.6 to equivariantly retract $K_{\Delta}$ to a complex $\widehat{K_{\Delta}}$ of dimension $M(L)=11$, concluding that $\textsc{vcd}(U(A_{\Delta}))=11$. Hence

### v2

@@ 7.1. A sufficient condition for $M(L)=M(V)$
Therefore $\overline{Q}$ and $\overline{Q^{\prime}}$ are both hugged. The only vertex which dominates $u_{1}$ and is contained in $\overline{Q}$ is $a_{2}$ – so the partition(s) which hug $\overline{Q}$ are based at $a_{2}$. In a hugging set for $\overline{Q}$, there must be a partition $\mathcal{P}$ which has side $P\subseteq\overline{Q}$ with $u_{2}\in P$. Since $\{a_{2}^{\pm}\}\subseteq\overline{Q^{\prime}}$, both sides of $\mathcal{P}$ must intersect $\overline{Q^{\prime}}$. Since $u_{2}\in Q^{\prime}\cap P$, we are forced to have $\overline{P}\cap Q^{\prime}=\emptyset$. Hence $P\supseteq\{u_{2},a_{3}^{\pm},b_{3}^{\pm}\}$.
We already said that the elements $\{u_{2}\},\{a_{3}^{\pm},b_{3}^{\pm}\}$ of $\mathcal{C}(a_{2})$ must be on different sides of $\mathcal{R}$. By the previous paragraph, both sides of $\mathcal{R}$ thus have non-empty intersection with $P$. But the elements $\{u_{1}\},\{a_{1}^{\pm},b_{1}^{\pm}\}$ of $\mathcal{C}(a_{2})$ are also on different sides of $\mathcal{R}$ – so both sides of $\mathcal{R}$ intersect $\overline{P}$. Hence $\mathcal{P}$ is incompatible with $\mathcal{R}$, which is a contradiction.
Hence, even though $\Delta$ violates Condition 2, it still satisfies the conclusion of Lemma 4.14. Since $\Delta$ is barbed, we can apply Corollary 5.6 to $U(A_{\Delta})$-equivariantly retract $K_{\Delta}$ to a complex $\widehat{K_{\Delta}}$ of dimension $M(L)=11$, concluding that $\textsc{vcd}(U(A_{\Delta}))=11$. Hence